Operations Report – February 8th



Crew 274 Operations Report 2/8/2023

SOL: 10

Name of person filing report: Alexis Lojek

Non-nominal systems: Toilet. Robotic Observatory.

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews. Robotic observatory is closed until further notice.


Spirit rover used: Yes

Hours: 208.5

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.5

Beginning charge: 100%

Ending charge: 76%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 220.7

Beginning charge: 100%

Ending charge: 80%

Currently charging: No

Perseverance rover used: Yes

Hours: 255.4

Beginning charge: 100%

Ending charge: 71%

Currently charging: No

General notes on rovers: Rovers Spirit and Opportunity were used for the morning EVA and plugged in when they returned, as it was a cloudless day and before noon. Curiosity and Perseverance were utilized for the afternoon EVA, but returned after noon, so were not plugged.

Summary of Hab operations: Toilet tank was emptied this morning and habitat is in a nominal condition.

WATER USE: 36 gallons as of 5:30pm

Water (static tank): 278 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged yesterday and are off chargers. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 1:01 – 1:32 pm.

WATER USE: 12.68 gallons

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: No.

Summary of ScienceDome operations: GHO Tyler Hines’ experiment is continuing to be monitored. Forty LDEF seeds placed in the GreenHab are continuing to be monitored as well.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The MDRS Robotic Observatory is inoperable until further notice due to a mechanical issue described in Sol 9’s Operations report.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site